# A quantum computing concept for 1-D elastic wave simulation

### v1

@@ 6 Conclusions and outlook
We presented a quantum computing concept for 1-D elastic wave propagation through heterogeneous media. Our approach involves (1) a finite-difference approximation of the wave equation, (2) a sparsity-preserving transformation to a Schrödinger equation, and (3) direct simulation of the Schrödinger equation on a gate-based quantum computer.
The implementation of our approach on an error-free quantum simulator, produces nearly exact results and serves as the foundation of wave field simulations with small numbers of grid points on the real quantum computer IBM Brisbane. While the simulation results qualitatively agree with the error-free version, they are contaminated by errors related to quantum decoherence and noise, which are typical issues for today’s quantum computers.
A continuous version of the discrete transformation to the Schrödinger equation enables the modification of our method based on finite differences, including the use of other spatial discretisation schemes, such as the spectral-element method. Assuming that error-corrected quantum chips will become available, an analogy between our method and coupled mass-spring systems suggests that our quantum computing approach may lead to wave field simulations that run exponentially faster than simulations on classical computers. This perspective motivates ongoing research on the extension of the 1-D elastic case to 3-D scenarios that include visco-elastic attenuation.
@@ Acknowledgements.
We would like to thank the other members of the ETH Seismology & Wave Physics Group for fruitful discussions and inspiration. All codes used in this work are available at https://github.com/malteschade/Quantum-Wave-Equation-Solver.
@@
Author contribution statement: theory and study conception: C. Bösch; implementation and execution on the quantum computer : M. Schade; quantum computing introduction: V. Hapla; analysis and interpretation of results: M. Schade, C. Bösch, V. Hapla, A. Fichtner; manuscript preparation: M. Schade, C. Bösch, V. Hapla, A. Fichtner. All authors review the results and approved the final version of the manuscript.
@@ Appendix A Continuous transformation
Here, we construct an invertible transformation $\mathcal{T}$ that transforms the continuous 1-D elastic wave equation to a continuous Schrödinger equation. We consider functions $a(x)$ that are differentiable at least once and satisfy either $a(0)=f(W)=0$, $a(0)=0$ and $\partial_{x}a(W)=0$, or $\partial_{x}a(0)=0$ and $a(W)=0$, which define three different Hilbert spaces $\mathbb{H}_{D,D}$, $\mathbb{H}_{D,N}$ or $\mathbb{H}_{N,D}$ referring to either two Dirichlet or at least one Dirichlet boundary condition. All three Hilbert spaces are equipped with the natural inner product
@@
for $B,B^{\prime}\in\{D,N\}$, and where *** denotes complex conjugation. We introduce the adjoint operator, $\mathcal{A}^{\dagger}$ of some operator $\mathcal{A}$, corresponding to the above inner product and satisfying
@@ Appendix B Complexity of sparse Hamiltonian simulation
To apply the Hamiltonian simulation approach from section 2.3, the quantum circuit needs to be properly initialised per (28). Today’s best sparse state initialisation methods need $\mathcal{O}(nS^{2}\log(S))$ classical operations, producing a circuit with $\mathcal{O}(nS)$ quantum gates, where $n$ is the qubit count and $S$ is the number of non-zero initial conditions (Gleinig and Hoefler, 2021). This cost can be amortised with relative ease if $S\ll 2^{n}$, which matches typical use cases.
To provide insight into the complexity of the time evolution itself using the currently best available methods, we first need to discuss the implementation of sparse matrices in general. Sparse matrices are typically implemented on QCs using oracles, which are functions that map one integer vector into another. A quantum oracle must be a unitary operation, and it can be shown that this requirement can be satisfied for any classical function by augmenting both its input and output with additional helper qubits appropriately. To represent a $d$-sparse matrix $\mathbf{X}$, two oracles are used: (1) one that takes a row index $i$ and column index $j$, and returns the value $\mathbf{X}(i,j)$, and (2) one that takes a row index $i$ and nonzero index $l\in[0,d)$, and returns the column index $j$ of the $l$-th nonzero element in the row $i$.
As described in section 2.3, the action of the time-evolution operator $e^{-i\mathbf{H}t}$ is carried out only approximately because the exact matrix exponential evaluation is prohibitively expensive. There are multiple algorithms available that realise this approximation, the most efficient so far being HS by qubitisation (Low and Chuang, 2019). The input $d$-sparse Hamiltonian $\mathbf{H}$ can be provided using the two oracles from above. For the time $t$ and error $\varepsilon$, the qubitisation algorithm needs $\mathcal{O}(dt||\mathbf{H}||_{\mathrm{max}}+\log(1/\varepsilon))$ queries to the oracles. Here, $||\mathbf{H}||_{\mathrm{max}}$ is the maximum absolute value of $\mathbf{H}$. The error $\varepsilon$ is the desired fidelity of the time-evolution operator, more specifically, the maximum admissible $||\mathbf{U}^{\prime}-\mathbf{U}||_{2}$, where $\mathbf{U}=e^{-it\mathbf{H}}$ denotes the exact time-evolution operator, and $\mathbf{U}^{\prime}$ denotes its approximation produced by the algorithm. While HS is an area of active research, the query complexity of this algorithm already reaches fundamental lower bounds in the $\mathcal{O}$-notation. In the presented 1-D case, $d=2$, as we have a 2-sparse Hamiltonian, and
@@
Hence, the query complexity is linear in $d$, $t$, $||\mathbf{H}||_{\mathrm{max}}$ and $1/\varepsilon$. Assuming that $\varepsilon$, $d$ and $\mathbf{H}_{\textrm{max}}$ do not depend on $n$, which clearly holds in our case here, the query complexity is constant in $n$, i.e., does not depend on the matrix size. It is usually assumed that the oracles themselves can be implemented with the polynomial runtime complexity, i.e., the circuit depth $D$ is in $\mathcal{O}(n^{k})$, implying the complexity of the whole Hamiltonian simulation is polynomial in $n$ as well. However, bounds on the runtime complexity of a query are an area of active research (Zhang and Yuan, 2023). A more technical, in-depth analysis of the sparse Hamiltonian simulation utilised for simulating coupled classical oscillators and proofs establishing exponential speedup against classical algorithms can be found in Babbush et al. (2023).
@@ Appendix C General Boundary Conditions
The derived Hamiltonian formulation based on the FD matrix (14) naturally results in Dirichlet and Neumann boundary conditions (BCs) on the right and left side of the domain, respectively. We now extend this approach to arbitrary BCs, with at least one Dirichlet BC to ensure negative definiteness of the stiffness matrix. To that end, we define the non-square first-order-accurate forward finite-difference matrix $\mathbf{D}$ as
@@
which we use in the Hamiltonian simulation. Note this formulation allows for $12^{n-1}-$ 1 grid points for $n$ qubits.

### v2

@@ 6 Conclusions and outlook
We presented a quantum computing concept for 1-D elastic wave propagation through heterogeneous media. Our approach involves (1) a finite-difference approximation of the wave equation, (2) a sparsity-preserving transformation to a Schrödinger equation, and (3) a direct simulation of the Schrödinger equation on a gate-based quantum computer.
The implementation of our approach on an error-free quantum simulator produces nearly exact results and serves as the foundation of wave field simulations with small numbers of grid points on the real quantum computer IBM Brisbane. While the simulation results qualitatively agree with the error-free version, they are contaminated by errors related to quantum decoherence and noise, which are typical issues for today’s quantum computers.
A continuous version of the discrete transformation to the Schrödinger equation enables the modification of our method based on finite differences, including the use of other spatial discretization schemes, such as the spectral-element method. Finally, assuming that error-corrected quantum chips will become available, we discuss that the best quantum algorithms can solve the elastic wave equation with exponential speedup over classical solvers.
With the FWI as an end goal, our ongoing research focuses on incorporating sources, measuring misfit functions, extending the formalism to higher space dimensions, including visco-elastic attenuation, and computing gradients.
@@ Acknowledgements.
We would like to thank the other members of the ETH Seismology & Wave Physics Group for fruitful discussions and inspiration. We would also like to acknowledge Shahpoor Moradi, Alexandre Souza, and one anonymous reviewer for insightful feedback. All codes used in this work are available at https://github.com/malteschade/Quantum-Wave-Equation-Solver.
@@
Author contribution statement: theory and study conception: C. Bösch; implementation and execution on the quantum computer: M. Schade; quantum computing introduction: V. Hapla; analysis and interpretation of results: M. Schade, C. Bösch, V. Hapla, A. Fichtner; manuscript preparation: M. Schade, C. Bösch, V. Hapla, A. Fichtner. All authors reviewed the results and approved the final version of the manuscript.
@@ Appendix A Continuous transformation
Here, we construct an invertible transformation $\mathcal{T}$ that transforms the continuous 1-D elastic wave equation into a continuous Schrödinger equation. We consider functions $a(x)$ that are differentiable at least once and satisfy either $a(0)=f(W)=0$, $a(0)=0$ and $\partial_{x}a(W)=0$, or $\partial_{x}a(0)=0$ and $a(W)=0$, which define three different Hilbert spaces $\mathbb{H}_{D,D}$, $\mathbb{H}_{D,N}$ or $\mathbb{H}_{N,D}$ referring to either two Dirichlet or at least one Dirichlet boundary condition. All three Hilbert spaces are equipped with the natural inner product
@@
for $B,B^{\prime}\in\{D,N\}$, and where * denotes complex conjugation. We introduce the adjoint operator, $\mathcal{A}^{\dagger}$ of some operator $\mathcal{A}$, corresponding to the above inner product and satisfying
@@ Appendix B Complexity of sparse Hamiltonian simulation
The first step of the stated HS approach is setting up the initial state according to (28). Today’s best sparse state initialisation methods need $\mathcal{O}(nS^{2}\log(S))$ classical operations, producing a circuit with $\mathcal{O}(nS)$ quantum gates, where $S$ is the number of non-zero initial conditions (Gleinig and Hoefler, 2021). This cost can be amortized with relative ease if $S\ll 2^{n}$, which matches typical use cases.
To provide insight into the complexity of the time evolution itself using the currently best available methods, we first need to discuss the implementation of sparse matrices. It is generally assumed that sparse matrices can efficiently be accessed using quantum oracles. A quantum oracle is a function solving a subroutine of an algorithm that has an efficient implementation, meaning it has a polynomial runtime complexity and hence its circuit depth $D$ is in $\mathcal{O}(n^{k})$. The existence of these oracles is generally assumed in quantum algorithmic research and allows for the analysis and improvement of subroutines in a larger algorithmic framework. A quantum oracle must be a unitary operation, and it can be shown that this requirement can be satisfied for any classical function by augmenting both its input and output with additional helper qubits appropriately. To represent a $d$-sparse matrix $\mathbf{X}$, two oracles are used: (1) one that takes a row index $i$ and column index $j$, and returns the value $\mathbf{X}(i,j)$, and (2) one that takes a row index $i$ and nonzero index $l\in[0,d)$, and returns the column index $j$ of the $l$-th nonzero element in the row $i$.
As described in section 2.3, the action of the time-evolution operator $e^{-i\mathbf{H}t}$ is carried out only approximately because the exact matrix exponential evaluation is prohibitively expensive. There are multiple algorithms available that realize this approximation, the most efficient so far being HS by qubitization (Low and Chuang, 2019). The input $d$-sparse Hamiltonian $\mathbf{H}$ can be provided using the two oracles from above. For the time $t$ and error $\varepsilon$, the qubitization algorithm needs $\mathcal{O}(dt||\mathbf{H}||_{\mathrm{max}}+\log(1/\varepsilon))$ queries to the oracles. Here, $||\mathbf{H}||_{\mathrm{max}}$ is the maximum absolute value of $\mathbf{H}$. The error $\varepsilon$ is the desired fidelity of the time-evolution operator, more specifically, the maximum admissible $||\mathbf{U}^{\prime}-\mathbf{U}||_{2}$, where $\mathbf{U}=e^{-it\mathbf{H}}$ denotes the exact time-evolution operator, and $\mathbf{U}^{\prime}$ denotes its approximation produced by the algorithm. While HS is an area of active research, the query complexity of this algorithm already reaches fundamental lower bounds in the $\mathcal{O}$-notation. In the presented 1-D case, $d=2$, as we have a 2-sparse Hamiltonian, and
@@
Hence, the query complexity is linear in $d$, $t$, $||\mathbf{H}||_{\mathrm{max}}$ and $1/\varepsilon$. Assuming that $\varepsilon$, $d$ and $\mathbf{H}_{\textrm{max}}$ do not depend on $n$, which clearly holds in our case, the complexity of the overall algorithm is polynomial in $n$ and logarithmic in $N$. This establishes the exponential speedup over classical solutions.
A more technical, in-depth analysis of the sparse HS utilized for simulating coupled classical oscillators and proofs establishing exponential speedup against classical algorithms can be found in Babbush et al. (2023) and a general complexity analysis for simulating $d$-sparse Hamiltonians in Low and Chuang (2019).
@@ Appendix C General Boundary Conditions
The derived Hamiltonian formulation based on the FD matrix (14) naturally results in Dirichlet and Neumann boundary conditions (BCs) on the right and left side of the domain, respectively. We now extend this approach to arbitrary BCs, with at least one Dirichlet BC to ensure the negative definiteness of the stiffness matrix. To that end, we define the non-square first-order-accurate forward finite-difference matrix $\mathbf{D}$ as
@@
which we use in the HS. Note this formulation allows for $12^{n-1}-$ 1 grid points for $n$ qubits.
Appendix D Quantum Circuit Implementation
Figure 3 displays the quantum circuit implemented in section 4 in the standard circuit model notation for one randomly chosen observable. While Figure 3(a) shows the circuit in a general and simplified form, Figure 3(b) shows the detailed circuit implementation for the circuit run on the QC IBM Brisbane. In both circuit models, we can observe the four distinct subroutines included in the full quantum algorithm: 1.) The state preparation of |$\ket{\psi(0)}$, 2.) the time evolution from |$\ket{\psi(0)}\rightarrow\ket{\psi(t)}$, 3.) the application of the observable to bring the state into the computational basis, and 4.) a full state measurement of all qubits in the computational basis.
After initializing the qubits in state |$\ket{0}$, a state preparation routine is employed. This sub-circuit prepares the initial wave field state from |$\ket{0}\rightarrow\ket{\psi(0)}$ (28). Generally, this state initialization can be done using sparse initialization routines (Gleinig and Hoefler (2021)), or simplified routines to initialize highly symmetric states with minimal circuit overhead. A variant of a simplified initialization routine is displayed in Figure 3(b). By using only three additional gates, it allows the initialization of a point-localized initial condition in the center of the simulated domain. We employ this simplified initialization routine in the simulation used for obtaining Figure 2.
As a next step, the time evolution circuit is implemented. This circuit evolves the state |$\ket{\psi(0)}$ to |$\ket{\psi(t)}$ (30). We construct the time evolution circuit from the Hamiltonian $H$ and the evolution time $t$ using a scaling and squaring method (Al-Mohy and Higham (2010)). We do not provide a detailed circuit representation of the time evolution sub-circuit as even for small problem sizes it contains hundreds of gates arranged in a non-trivial layout.
Finally, we implement an observable according to the QST requirements. As we need to evaluate |$\ket{\psi(t)}$ with respect to $2^{n}$ quantum observables for every time step (see section 4), the circuits displayed in Figure 3 describe only one of the many necessary observable sub-circuit layouts evaluated on the QC. As each circuit evaluation only gives one measurement sample, every circuit layout needs to be evaluated multiple times to obtain statistically accurate results. To implement an observable given as a Pauli matrix on a QC it is necessary to rotate all qubits into the computational basis or $Z$-basis. The rotation directions are defined by the respective Pauli matrix components for each qubit. Figure 3(b) shows an implementation of the Pauli observable $XZXZ$, which equates to implementing Hadamard gates ($H$) for every $X$ component, and no additional gates for every $Z$ component. This sub-circuit achieves a rotation of all qubits into the $Z$-basis, where they can be measured.
As a last step, all qubits are measured and the results are subsequently used in the QST process.
The full circuit implementation is available at https://github.com/malteschade/Quantum-Wave-Equation-Solver.
(a) Simplified Circuit.
(b) Expanded Circuit
Figure 3: Circuit Representation of the wave simulation algorithm for a 7 grid point problem subject to the XZXZ Pauli observable. The circuits show four distinct algorithmic stages in differing levels of detail: 1.) The state preparation of |$\ket{\psi(0)}$, 2.) the time evolution from |$\ket{\psi(0)}\rightarrow\ket{\psi(t)}$, 3.) the application of the observable rotating the state into the computational basis, and 4.) a full state measurement of all qubits in the computational basis.